# The deep multi-FBSDE method: a robust deep learning method for coupled FBSDEs

### v2

@@ 4 Numerical experiments
We present four numerical experiments, two with an associated stochastic optimal control problem, and two without such a connection, all with analytic or semi-analytic reference solutions so that we can analyze the error. The first problem is a controlled Brownian motion in two dimensions, the second problem is a six-dimensional LQ-problem, and the last two examples are FBSDEs related to two-dimensional advection diffusion reaction PDEs, one without a reaction term and one with a linear one. We demonstrate the performance of both the deep BSDE and the deep multi-FBSDE methods.
In the literature, it is most common to demonstrate deep BSDE-type methods on very high-dimensional problems. However, these problems are most often very symmetric, with solutions being permutation-invariant in its space variables. It is highly questionable how important these examples are for applications and also if they fairly represent more general problems of the same dimensions. We have instead chosen difficult low-dimensional problems for which the classical deep BSDE methods fail, rather than high-dimensional problems.
In the implementation of the deep BSDE and both Phase I–II of the deep multi-FBSDE methods, we optimize $\zeta_{0},\dots,\zeta_{N-1}$ in a function class defined by ReLU feed-forward neural networks with 3 hidden layers with 20 nodes in each layer. The initial value $y_{0}$ in the deep BSDE method and in Phase I of the deep multi-FBSDE method is represented by one single trainable parameter. Training data consist of $2^{20}$ independent realizations of the $d$-dimensional Wiener increment vectors $\Delta W_{0},\ldots,\Delta W_{N-1}$, each element being i.i.d., and normally distributed with mean zero and variance $h$. The training data are reused in $10$ epochs. Training is initiated by randomly sampling parameters and proceeds with the classical Adam optimizer with batches of size $2^{12}$.
@@ 4.1 FBSDEs stemming from stochastic optimal control problems
The value function associated with the control problem is formally defined as
@@ 4.1.1 Controlled Brownian motions with general terminal cost
together with a flexibly chosen terminal cost function $g\colon\mathbb{R}^{d}\to\mathbb{R}$. Straightforward calculations yield the optimal control for this problem $u^{*}=-\frac{1}{r}\mathrm{D}_{x}v=-\frac{1}{r}Z$, from which we obtain the FBSDE
@@
Using the Cole–Hopf transformation, the solution to the above FBSDE is for $t\in[0,T]$, recovered as
$Y_{t}=-\frac{1}{r\sigma^{2}}\log\bigg{(}E\big{[}\text{e}^{-r\sigma^{2}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\big{]}\bigg{)},$ (16)
@@
The FBSDE (15) can then be estimated using Monte Carlo sampling. It is important to note that, in approximating $(Y_{t}(\omega))_{t\in[0,T]}$, nested expectations along the trajectory of $(X_{t}(\omega))_{t\in[0,T]}$ are required. Moreover, since $X$ is coupled with $Z$ in the drift component, straightforward application of this strategy for approximating $Y$ at times other than $t=0$ (which is independent of $X$’s trajectory) is challenging. However, if our approximation of $(X,Y,Z)$ is accurate, it should follow that our approximation of $Y$ aligns with the Monte Carlo approximation of $Y$ using (16), along the simulated trajectories of $X$.
@@
We set $T=0.5$, $d=\ell=k=2$, $r=1$, and $\sigma=0.25$. The initial state is set at $x_{0}=(-0.1,0.1)^{\top}$. We employ a terminal cost function $g(x)=-|x_{1}-x_{2}|$, where $x=(x_{1},x_{2})^{\top}$. This configuration aims to control the two components of the Brownian motion such that they diverge as much as possible by the time $t=T$, while minimizing the use of control force. The discretization of the problem is set with $N=40$.
@@
While the deep FBSDE method fails to converge to the true solution, the reason behind this discrepancy remains unclear. To further explore this discrepancy, we illustrate in Figure 7 the sample mean of the forward SDE along with a representative path of the backward SDE, comparing these with a reference solution computed using the Cole–Hopf transformation and Monte Carlo sampling.
@@
Intuitively, the optimal strategy should be to direct $X^{1}$ downwards and $X^{2}$ upwards. This behavior is observed in the deep multi-FBSDE method, whereas the deep FBSDE method results in the opposite: $X^{1}$ is driven upwards while $X^{2}$ is driven downwards. Interestingly, around $t=0.18$, the first component of $X$ surpasses the second component, and from this point forward, the control actions appear nearly optimal given the state configuration. This assertion is partially supported by the trajectory of the approximate $Y$, which closely follows its analytic counterpart. It is important to understand that the accuracy of the analytic $Y$ is conditioned on the trajectory of $X$ and is subject to Monte Carlo error. It is even crucial to note that the behavior described above does not stem from a local minimum encountered during the training process. Employing a similar methodology to that used in previous sections, where the mean-squared error is calculated for various initial values of the discretized BSDE, we confirm that the observed solution represents a global minimum. Therefore, this outcome is consistent with what should be expected a priori, given the problem setup and model parameters.
@@ 4.1.2 Linear Quadratic Gaussian control problems
The above, together with (12), defines the state equation and cost functional for a LQ control problem. The optimal feedback control, minimizing the Hamiltonian, $\inf_{u\in U}{\langle\text{D}xV,Bu\rangle+\langle R_{u}u,u\rangle}$, is given by
$u^{*}_{t}=-\frac{1}{2}R_{u}^{-1}B^{T}\text{D}_{x}v(t,X_{t}).$ (18)
@@
$\begin{dcases}\dot{P}(t)-A^{T}P(t)-P(t)A-P(t)BR_{u}^{-1}B^{T}P(t)+R_{x}=\mathbf{0}_{d\times d},\\ \dot{Q}(t)+2P(t)AC-A^{T}Q(t)-P(t)BR_{u}^{-1}B^{T}Q(t)=\mathbf{0}_{d},\\ \dot{R}(t)+\text{Tr}\big{(}\sigma\sigma^{\top}P(t)\big{)}+Q(t)^{\top}AC-\frac{1}{4}Q(t)^{\top}BR_{u}^{-1}B^{\top}Q(t)=0,\quad t\in[0,T],\\ P(T)=G;\quad Q(T)=\mathbf{0}_{d};\quad R(T)=0.\end{dcases}$
@@
For the deep multi-FBSDE method, we set $K=3$ and define $\psi_{1}(t,x,y,z)=A(C-x)$, $\psi_{2}(t,x,y,z)=-A(C-x)$, and $\psi_{3}(t,x,y,z)=0$. With this configuration, we incorporate both the doubling effect (through $\psi_{2}$) and the removal of the mean-reverting term (through $\psi_{1}$). Furthermore, since $\psi_{3}$ is identically zero, the original control problem is also explicitly included in the training, which appears to have a regularizing effect on the numerical results.
The matrices used for the state equation are given by
@@
$\sigma$ $=\text{diag}([0.2,1,0.2,1,0.2,1]),\quad x_{0}=(0.1,0.1,0.1,0.1,0.1,0.1)^{\top},\quad T=0.5.$
The penalty matrices of the control problem are given by
@@
Figure 8 clearly demonstrates that our method yields accurate approximations for $(X,Y,Z)$, both pathwise and in distribution. We refrain from comparing these results with those obtained via the deep BSDE method, as the latter produces completely incorrect approximations for this problem. The approximate initial condition obtained by the deep BSDE method is around 28.44, whereas the true value is approximately 9 (see Figure 1). In [30], the authors use the deep BSDE method to approximate the solution of a BSDE derived from the stochastic maximum principle for the specific problem discussed above. However, this problem is significantly less complex because, at each time $t$, the control process in the associated BSDE is constant in the state space. Consequently, it remains unclear to what extent this approach would generalize to more complex problems where the control process is non-constant.
@@
Figure 8: Average of solutions and a single solution path compared to their semi-analytic counterparts for the LQ-control problem from Section 4.1.2. The shaded area represents an empirical credible interval for $Y$, defined as the area between the 5:th and the 95:th percentiles at each time point. We do not include credible intervals for $X$ and $Z$ in this figure to facilitate visualization. For $X$ and $Z$, we see one realization of each of the six components.
@@ 4.2 FBSDEs from advection diffusion reaction PDEs
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}\beta Z_{t}\text{d}s+\sqrt{2\alpha}W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\big{(}\beta\|Z_{s}\|^{2}-\gamma Y_{s}\big{)}\text{d}s-\sqrt{2a}\int_{t}^{T}\langle Z_{s},\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (23)
@@
$\begin{dcases}\frac{\partial u}{\partial t}(t,x)+\alpha\Delta u-cu\log{u}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ u(T,x)=\exp\Bigg{(}\frac{2\beta}{\alpha}g(x)\Bigg{)},&x\in\mathbb{R}^{d}.\end{dcases}$ (24)
@@
$Y_{t}=-\frac{2\beta}{\alpha}\log{\bigg{(}E\Big{[}\text{e}^{-\frac{\alpha}{2\beta}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\Big{]}\bigg{)}},$ (25)
@@
Figure 10 displays the approximate initial condition during training for both the deep BSDE and deep multi-FBSDE methods.
@@
The figure clearly illustrates that the deep BSDE method yields inaccurate approximations, whereas the deep multi-FBSDE method produces highly accurate results. For the second experiment, we explore the accuracy of the deep BSDE method and the deep multi-FBSDE method for different values of $\beta$. In Figure 11, we see that for $\beta>0.25$, the deep BSDE fails to provide accurate approximations. The deep multi-FBSDE method provides accurate approximations for all $\beta$ tested.
Figure 11: Approximate initial values for different values of $\beta$. To not distort visibility of the plot we have used the maximum of $-1$ and the approximate initial conditions ($Y_{0}$ approximated with the deep BSDE method for $\beta=10$, $\beta=1.25$ and $\beta=1$ is $-1.04$, $-2.9$ and $-4.9$, respectively).
@@ 5 Conclusions and potential future research directions
In this paper, we introduced a two-phase deep-learning-based method, the deep multi-FBSDE method, for approximating coupled forward–backward stochastic differential equations (FBSDEs). Our approach targets scenarios where the standard deep BSDE method [17] fails to converge. The main characteristic of our method is that its loss function is composed of the sum of multiple deep BSDE loss functions, each originating to a FBSDE satisfying the same partial differential equation (PDE) as the original FBSDE, but with transformed drift and driver. The method was demonstrated on two stochastic control problems and two FBSDEs related to advection diffusion reaction PDEs, unrelated to stochastic control. For neither of the problems, the deep BSDE method worked, but the deep multi-FBSDE method did. Overall, our results suggest that focusing on a family of equivalent FBSDEs offers a promising path to overcome convergence challenges in deep BSDE methods, thus broadening the scope of problems in finance, control theory, and related areas where these methods can be applied effectively.
Although the method seems to work very well on challenging FBSDEs, we still do not understand the reason for this and why the deep BSDE method fails. There is thus a need to both understand theoretically for what problems the deep BSDE does not work and why, and why the deep multi-FBSDE method does work for these problems. From a practical perspective, the choice of $K$ and $\psi_{1},\dots,\psi_{K}$ in our algorithm is currently somewhat ad hoc. While this works for moderately complex problems, a more systematic approach to selecting $K$ and $\psi_{1},\dots,\psi_{K}$ would be especially beneficial when dealing with high-dimensional systems or models where it is difficult to anticipate the impact of different choices. Developing such a framework constitutes a natural next step toward enhancing both the reliability and the applicability of the proposed method. Furthermore, extending the methodology to broader classes of PDE-driven problems and exploring theoretical guarantees for large-scale or high-dimensional settings would deepen our understanding of its robustness and versatility.

### v3

@@ 4 Numerical experiments
We present four numerical experiments, two with an associated stochastic optimal control problem, and two without such a connection, all with analytic or semi-analytic reference solutions so that we can analyze the error. The first problem is a controlled Brownian motion in two dimensions, the second problem is a six-dimensional LQ-problem, and the last two examples are FBSDEs related to two-dimensional advection diffusion reaction PDEs, one without a reaction term and one with a linear one. We demonstrate the performance of both the deep FBSDE and the deep multi-FBSDE methods.
In the literature, it is most common to demonstrate deep BSDE-type methods on very high-dimensional problems. However, these problems are most often very symmetric, with solutions being permutation-invariant in its space variables. It is highly questionable how important these examples are for applications and also if they fairly represent more general problems of the same dimensions. We have instead chosen difficult low-dimensional problems for which the classical deep FBSDE methods fail, rather than high-dimensional problems.
In the implementation of the deep FBSDE and both Phase I–II of the deep multi-FBSDE methods, we optimize $\zeta_{0},\dots,\zeta_{N-1}$ in a function class defined by ReLU feed-forward neural networks with 3 hidden layers with 20 nodes in each layer. The initial value $y_{0}$ in the deep FBSDE method and in Phase I of the deep multi-FBSDE method is represented by one single trainable parameter. Training data consist of $2^{20}$ independent realizations of the $d$-dimensional Wiener increment vectors $\Delta W_{0},\ldots,\Delta W_{N-1}$, each element being i.i.d., and normally distributed with mean zero and variance $h$. The training data are reused in $10$ epochs. Training is initiated by randomly sampling parameters and proceeds with the classical Adam optimizer with batches of size $2^{12}$.
@@ 4.1 FBSDEs stemming from stochastic optimal control problems
The controls $u$ belong to a suitable subset of adapted stochastic processes, and we refrain from details The value function associated with the control problem is formally defined as
@@ 4.1.1 Controlled Brownian motions with general terminal cost
together with a flexibly chosen terminal cost function $g\colon\mathbb{R}^{d}\to\mathbb{R}$. Straightforward calculations yield the optimal control $u^{*}=-\frac{1}{r}\mathrm{D}_{x}v=-\frac{1}{r}Z$, from which we obtain the FBSDE
@@
Using the Cole–Hopf transformation, the unique $Y$-component of the solution to the above FBSDE is
$Y_{t}=-r\sigma^{2}\log\bigg{(}\mathbb{E}\big{[}\text{e}^{-\frac{1}{r\sigma^{2}}g(X_{t}+\sigma\sqrt{T-t}\,\xi)}\big{|}X_{t}\big{]}\bigg{)},\quad t\in[0,T],$ (16)
@@
The FBSDE (15) can then be approximated using Monte Carlo sampling. It is important to note that, to approximate $(Y_{t}(\omega))_{t\in[0,T]}$, nested expectations are required along the trajectory of $(X_{t}(\omega))_{t\in[0,T]}$. Moreover, since $X$ is coupled with $Z$ in the drift component, a straightforward application of this strategy to approximate $Y$ at times other than $t=0$ (which is independent of the trajectory of $X$) is challenging. However, if our approximation of $(X,Y,Z)$ is accurate, it should follow that our approximation of $Y$ aligns with the Monte Carlo approximation of $Y$ using (16), along the simulated trajectories of $X$.
@@
We set $T=0.5$, $d=\ell=k=2$, $r=1$, and $\sigma=0.25$. The initial state is set at $x_{0}=(-0.1,0.1)^{\top}$. We employ a terminal cost function $g(x)=-|x_{1}-x_{2}|$, where $x=(x_{1},x_{2})^{\top}$. This configuration aims to control the two components of the Brownian motion such that they diverge as much as possible by time $t=T$, while minimizing the use of control force. The discretization of the problem is set with $N=40$.
@@
The deep FBSDE method fails to converge to the true solution. Although we do not have theoretical understanding for the reason, to gain some partial insight, we illustrate in Figure 7 the sample mean and a representative path of $X$ and $Y$, for the two numerical methods and the reference solution.
@@
We see that the optimal strategy pushes $X^{1}$ to lower values and $X^{2}$ to higher values, which is intuitive since $X_{0}^{1}<X_{0}^{2}$ and the task is to maximize $|X_{T}^{1}-X_{T}^{2}|$ with low control cost. The deep FBSDE method, on the other hand, causes $X^{1}$ and $X^{2}$ to instead cross each other, and naturally uses a more expensive control for the task, reading the initially high values of the $Y$-process. After the crossing, the control seems to resemble the optimal control well, which can be read from the resulting $Y$-process being close to the reference. A similar MSE-analysis as in Figure 1 (omitted) shows that the problem with the deep FBSDE solution is due to a bad global minimum.
@@ 4.1.2 Linear Quadratic Gaussian control problems
The above, together with (12), defines the state equation and cost functional for a LQ control problem. The optimal feedback control, minimizing the Hamiltonian, $\inf_{u\in U}{\langle\text{D}_{x}V,Bu\rangle+\langle R_{u}u,u\rangle}$, is given by
$u^{*}_{t}=-\frac{1}{2}R_{u}^{-1}B^{\top}\text{D}_{x}v(t,X_{t}).$ (18)
@@
$\begin{dcases}\dot{P}(t)-A^{\top}P(t)-P(t)A-P(t)BR_{u}^{-1}B^{\top}P(t)+R_{x}=\mathbf{0}_{d\times d},\\ \dot{Q}(t)+2P(t)AC-A^{\top}Q(t)-P(t)BR_{u}^{-1}B^{\top}Q(t)=\mathbf{0}_{d},\\ \dot{R}(t)+\text{Tr}\big{(}\sigma\sigma^{\top}P(t)\big{)}+Q(t)^{\top}AC-\frac{1}{4}Q(t)^{\top}BR_{u}^{-1}B^{\top}Q(t)=0,\quad t\in[0,T],\\ P(T)=G;\quad Q(T)=\mathbf{0}_{d};\quad R(T)=0.\end{dcases}$
@@
For the deep multi-FBSDE method, we set $K=3$ and define $\psi_{1}(t,x,y,z)=A(C-x)$, $\psi_{2}(t,x,y,z)=-A(C-x)$, and $\psi_{3}(t,x,y,z)=0$. With this configuration, we incorporate both the doubling effect (through $\psi_{2}$) and the removal of the mean-reverting term (through $\psi_{1}$). Furthermore, since $\psi_{3}$ is identically zero, the original control problem is also explicitly included in the training.
The matrices used for the state equation are the same, up to $\sigma$, $R_{x}$ and $G$, as in [1, 33]. In fact $R_{x}=\mathrm{diag}(5,1,5,1,5)$ is $G=\mathrm{diag}(1,5,1,5,1)$ in the implementation of [1], but reads wrongly $R_{x}=\mathrm{diag}(25,1,25,1,25)$ and $G=\mathrm{diag}(1,25,1,25,1)$ in the presentation of that paper). Our setting reads
@@
$\sigma$ $=\text{diag}([0.2,1,0.2,1,0.2,1]),\quad x_{0}=(0.1,0.1,0.1,0.1,0.1,0.1)^{\top},\quad T=0.5,$
and
@@
In [33], the authors use the deep FBSDE method on their control problem, to approximate the solution of a BSDE derived from the stochastic maximum principle. However, this problem is structurally less complex because, at each time $t$, the control process in the associated BSDE is constant in the state space. It remains to understand to what extent this approach would generalize to more complex problems where the control process is non-constant.
@@
Figure 8 clearly demonstrates that our method yields accurate approximations for $(X,Y,Z)$, both pathwise and in distribution. We refrain from comparing these results with those obtained via the deep FBSDE method, as the latter produces completely incorrect approximations for this problem. The approximate initial condition obtained by the deep FBSDE method is around 28.44, whereas the true value is approximately 9 (see Figure 1).
@@
Figure 8: Empirical mean of solutions (left) and a single solution trajectory (right) compared to the reference solution for the LQ-control problem from Section 4.1.2. The shaded area represents an empirical credible interval for $Y$, defined as the area between the 5:th and the 95:th percentiles at each time point. We do not include credible intervals for $X$ and $Z$ in this figure to facilitate visualization. For $X$ and $Z$, we see one realization of each of the six components.
@@ 4.2 FBSDEs from advection diffusion reaction PDEs
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}\beta Z_{t}\text{d}s+\sqrt{2\alpha}W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\big{(}\beta\|Z_{s}\|^{2}-\gamma Y_{s}\big{)}\text{d}s-\sqrt{2\alpha}\int_{t}^{T}\langle Z_{s},\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (23)
@@
$\begin{dcases}\frac{\partial u}{\partial t}(t,x)+\alpha\Delta u-\gamma u\log{u}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ u(T,x)=\exp\Bigg{(}\frac{2\beta}{\alpha}g(x)\Bigg{)},&x\in\mathbb{R}^{d}.\end{dcases}$ (24)
@@
$Y_{t}=-\frac{\alpha}{2\beta}\log{\bigg{(}E\Big{[}\text{e}^{-\frac{2\beta}{\alpha}g(X_{t}+\sigma\sqrt{T-t}\,\xi)}\big{|}X_{t}\Big{]}\bigg{)}},$ (25)
@@
Figure 10 displays the approximate initial condition during training for both the deep FBSDE and deep multi-FBSDE methods.
@@
The figure clearly illustrates that the deep FBSDE method yields inaccurate approximations, whereas the deep multi-FBSDE method produces highly accurate results. For the second experiment, we explore the accuracy of the deep FBSDE method and the deep multi-FBSDE method for different values of $\beta$. In Figure 11, we see that for $\beta>0.25$, the deep FBSDE fails to provide accurate approximations. The deep multi-FBSDE method provides accurate approximations for all $\beta$ tested.
Figure 11: Approximate initial values for different values of $\beta$. Values of $Y_{0}$ smaller than $-1$ have been set to $-1$ for the sake of visibility.
@@ 5 Conclusions and potential future research directions
In this paper, we introduced a two-phase deep-learning-based method, the deep multi-FBSDE method, for approximating coupled forward–backward stochastic differential equations (FBSDEs). Our approach focuses on scenarios where the standard deep FBSDE method [20] does not converge. The main characteristic of our method is that its loss function is composed of the sum of multiple deep FBSDE loss functions, each originating from a FBSDE satisfying the same partial differential equation (PDE) as the original FBSDE, but with transformed drift and driver. The method was demonstrated on two stochastic control problems and two FBSDEs related to advection diffusion reaction PDEs, unrelated to stochastic control. For neither of the problems, the deep FBSDE method worked, but the deep multi-FBSDE method did. Overall, our results suggest that focusing on a family of equivalent FBSDEs offers a promising path to overcome convergence challenges in deep FBSDE methods, thus broadening the scope of problems in finance, control theory, and related areas where these methods can be applied effectively.
Although the method seems to work very well on challenging FBSDEs, we still do not understand the reason for this and why the deep FBSDE method fails. There is thus a need to both understand theoretically for what problems the deep FBSDE does not work and why, and why the deep multi-FBSDE method does work for these problems. From a practical perspective, the choice of $K$ and $\psi_{1},\dots,\psi_{K}$ in our algorithm is currently somewhat ad hoc. While this works for moderately complex problems, a more systematic approach to selecting $K$ and $\psi_{1},\dots,\psi_{K}$ would be especially beneficial when dealing with high-dimensional systems or models where it is difficult to anticipate the impact of different choices. Developing such a framework constitutes a natural next step toward enhancing both the reliability and the applicability of the proposed method. Furthermore, extending the methodology to broader classes of PDE-driven problems and exploring theoretical guarantees for large-scale or high-dimensional settings would deepen our understanding of its robustness and versatility.